OpenStudy (anonymous):

4. 2(x + 5) > 8x – 8 = 5.–2(x – 3) ≥ 5 – (x + 3) Please help and show me how to solve these inequalities.. i forget what to do first. thank you!!!!! :)

OpenStudy (anonymous):

4. 2x+10>8x-8 .......18>6x........x<3 distribute 2 first 5. -2x+6 >=5-x-3...-2x+6>=2-x......4>=x distribute -2 for left side and distribute - for right side and solve